### v5

# RSNet: A Light Framework for The Detection of SAR Ship Detection

Abstract
Recent advancements in synthetic aperture radar (SAR) ship detection using deep learning have significantly improved accuracy and speed, yet effectively detecting small objects in complex backgrounds with fewer parameters remains a challenge. This letter introduces RSNet, a lightweight framework constructed to enhance ship detection in SAR imagery. To ensure accuracy with fewer parameters, we proposed Waveletpool-ContextGuided (WCG) as its backbone, guiding global context understanding through multi-scale wavelet features for effective detection in complex scenes. Additionally, Waveletpool-StarFusion (WSF) is introduced as the neck, employing a residual wavelet element-wise multiplication structure to achieve higher dimensional nonlinear features without increasing network width. The Lightweight-Shared (LS) module is designed as detect components to achieve efficient detection through lightweight shared convolutional structure and multi-format compatibility. Experiments on the SAR Ship Detection Dataset (SSDD) and High-Resolution SAR Image Dataset (HRSID) demonstrate that RSNet achieves a strong balance between lightweight design and detection performance, surpassing many state-of-the-art detectors, reaching 72.5% and 67.6% in $\mathbf{mAP_{.50:.95}}$ respectively with 1.49M parameters. Our code will be released soon.
Index Terms:
synthetic aperture radar (SAR), target detection, lightweight model, multiscale feature fusion
I Introduction
Synthetic Aperture Radar (SAR) imagery is a vital tool in remote sensing [1], offering high-resolution images under all weather and lighting conditions. By using microwave signals that can penetrate atmospheric obstacles like clouds and rain, SAR provides detailed insights into terrain [2], vegetation [3], and man-made structures [4]. Its applications span geological exploration [5], disaster monitoring [6], agriculture [7], and urban planning. Notably, SAR excels in ocean monitoring [8], detecting subtle surface deformations, and generating digital elevation models, making it indispensable for modern earth observation.
With SAR image data scale and complexity increasing, traditional image processing methods are becoming insufficient [9]. Recent advances in convolutional neural networks (CNNs) and Transformer architectures have significantly improved target detection [10], [11]. These deep learning methods are generally categorized into candidate region-based and regression-based approaches, with the latter excelling in real-time detection by directly regressing targets at multiple scales. For SAR object detection, one-stage methods strike a promising balance between speed and accuracy; however, due to the computational and energy limitations of airborne platforms, a lightweight model framework is essential for practical deployment.
Building on the established model framework, several notable improvements have been proposed. For example, Tian et al. [12] introduced LFer-Net, which integrates SPDConv and InceptionDWConv for high accuracy and efficiency. Zhou et al. [13] proposed HRLE-SARDet, leveraging hybrid representation learning to improve speed-accuracy balance. Chang et al. [14] presented MLSDNet, utilizing Adaptive Scale Distribution Attention (ASA) to enhance generalization and detection precision. Feng et al. [15] developed LPEDet, an anchor-free SAR ship detection model with a novel position-enhanced attention strategy. Yang et al. [16] designed a lightweight network featuring bidirectional pyramidal structures and soft quantization to minimize accuracy loss and background interference.
While advancements have shown the viability of lightweight models for SAR detection, substantial improvement opportunities remain. SAR images, unlike optical images, are prone to speckle noise, which reduces detail resolution [1], and face increased interpretation complexity due to geometric distortions. Furthermore, the processing of grayscale SAR data lacks intuitive color information. These factors hinder the balance between accuracy and detection speed, particularly regarding the crucial metric $\mathrm{mAP_{.50:.95}}$ for optimizing detection performance.
To tackle challenges in remote sensing image detection, we present RSNet (Restar the Sky Network), an optimized network based on YOLOv8, featuring improvements to its backbone, neck, and head. The backbone incorporates the Waveletpool-ContextGuided (WCG) structure, while the head employs the Waveletpool-StarFusion (WSF) network. Additionally, a Lightweight-Shared (LS) module in the detection head reduces the model’s parameter load.
The primary contributions of this article are as follows:
  • We developed the WCG architecture, enhancing detection accuracy through anti-aliasing and contextual integration, achieving superior performance with fewer parameters.
  • The WSF network is constructed by combining wavelet upsampling pooling and element-wise multiplication through the residual structure, significantly reducing parameters while maintaining computational efficiency.
  • We introduced the LS module using the group normalization strategy to further minimize the parameter burden of the detection head, thereby enhancing the overall efficiency of RSNet.
  • Experiments conducted on the SSDD and the HRSID show that the proposed RSNet strikes an improved balance between lightweight design and detection performance.
II PROPOSED METHOD
Figure 1: Overall flow of the RSNet.
II-A Overall Architecture
This letter introduces RSNet (Restar the Sky Network), a lightweight network for ship detection in SAR images, structured into backbone, neck, and detect components as illustrated in Fig. 1. The backbone processes the input image through the first three layers to produce feature maps, which is then enhanced by multiple scales of wavelet pooling to generate multi-scale feature maps. The ContextGuided block further enriches feature extraction by integrating contextual information. In the neck, the WSF module employs wavelet upsampling to restore spatial resolution and fuse features. The lightweight star module combines an element-wise multiplication structure with skip connections for deeper feature extraction, enhancing small object recognition capability. Finally, the LS Detector module connects the features to output the bounding boxes, class labels, and confidence scores of the detected objects.
II-B Waveletpool-ContextGuided (WCG)
In complex backgrounds, including land structures, clutter, and coastlines, current SAR ship detectors face challenges in accurately locating ships. To achieve better performance with fewer parameters, this letter proposes WCG.
The WCG network introduces a wavelet pooling mechanism that efficiently extracts multi-scale spatial features, reducing computational complexity while enhancing sensitivity to subtle details in complex backgrounds. This allows the network to focus on potential ship targets in noisy environments. Additionally, the context guided convolutional module integrates surrounding environmental information, significantly improving detection of small targets, especially in high-noise conditions. By leveraging contextual information, the model better distinguishes real ship targets from background noise, enhancing localization accuracy. Below are specific descriptions of wavelet pooling and context guided convolution.
Figure 2: Flowchart of WSF (Neck of RSNet). The feature map from the backbone network is processed by a residual structure comprising four submodules: WaveletUnPool, Concat, and Star. The network structure of the Star module is also illustrated.
Figure 3: The structure of LS Detector.
II-B1 Wavelet Pooling
In the WCG, the WaveletPool module plays a critical role, designed to effectively extract multi-scale features from the input feature map through wavelet transforms. The output $Y_{pool}$ of this module is defined by the following formula:
$Y_{{pool}}={Conv}_{2{D}}(Y_{{C2f}})\in{R}^{B\times C_{{out}}\times H\times W},$ (1)
where we define the output feature map from C2f as $Y_{{C2f}}$. $B$ is the batch size. $H$ and $W$ are the dimensions of the downsampled output. $C_{{out}}$ is the number of output channels of the WaveletPool block. The filter set ${F}$ used in the WaveletPool operation consists of four different convolutional kernels defined as:
$F=\{F_{LL},F_{LH},F_{HL},F_{HH}\},$ (2)
$F_{LL}=\begin{bmatrix}0.5&0.5\\ 0.5&0.5\end{bmatrix},F_{LH}=\begin{bmatrix}-0.5&-0.5\\ 0.5&0.5\end{bmatrix},$
$F_{HL}=\begin{bmatrix}-0.5&0.5\\ -0.5&0.5\end{bmatrix},F_{HH}=\begin{bmatrix}0.5&-0.5\\ -0.5&0.5\end{bmatrix},$
where $F_{LL}$ is equivalent to average pooling. The WaveletPool operation can be expressed as:
$Y_{{pool}}=\sum_{j=1}^{4}F_{j}*Y_{{C2f}},$ (3)
where * denotes the convolution operation. This operation extracts multi-scale features from the input feature map by applying each filter $F_{j}$ and aggregating the results across the channel dimension.
II-B2 Context Guided
The ContextGuided block takes the input feature map and performs a series of operations, defined as:
$Y_{CGB}={CGB}(Y_{pool}).$ (4)
The processing steps within the ContextGuided block can be expressed as follows:
$X_{1}={Conv}_{1\times 1}(Y_{{pool}})\in{R}^{B\times n\times H\times W},n=\frac{C_{{out}}}{2},$ (5)
where $n$ represents the number of the input channels. Then a depthwise convolution is employed to capture local features $L$, and another depthwise convolution with a dilation rate $d$ is applied to capture the surrounding features $S$:
$L={Conv}_{3\times 3}(X_{1})\in{R}^{B\times n\times H\times W},$ (6)
$S={Conv}_{3\times 3}(X_{1})\quad\text{with dilation rate}\;d.$ (7)
The $L$ and $S$ are concatenated, normalized, and activated, generating the global context feature $F_{glo}$ through the global feature extraction module:
$J=BatchNorm({Concat}(L,S))\in{R}^{B\times 2n\times H\times W},$ (8)
$G=F_{\text{glo}}(J)\in{R}^{B\times C_{\text{out}}\times H\times W}.$ (9)
Finally, the output of the Context Guided Block is computed by adding the input to the global context refinement output:
$Y_{{CGB}}=Y_{{pool}}+G.$ (10)
II-C Waveletpool-StarFusion (WSF)
Although the WCG module reduces parameters, this letter proposes WSF to better balance parameters and model accuracy as described in Fig. 2. WaveletUnPool and Star module enhance SAR ship detection by restoring spatial resolution and retaining rich features for improved contour recognition. Star module utilizes depthwise separable convolutions and element-wise multiplication operations for real-time applicability while effectively mitigating noise. Together, they significantly enhance performance and efficiency in ship detection.
II-C1 WaveletUnPool
WaveletUnPool, in contrast, aims to restore the spatial resolution of the downsampled feature map, effectively reconstructing its original dimensions. It utilizes transpose convolution with a stride of 2, employing the same wavelet filters as described in formula 2 to upsample the feature map back to a higher resolution.
II-C2 Star
Element-wise multiplication structure inside StarNet [17] which is an advanced deep learning architecture that enhances model representation. By fusing features from different subspaces with element-wise multiplication at each layer, it expands implicit feature dimensions without significantly increasing parameters. Fig. 2 illustrates this process in the neck section, showing the initial concatenation of the feature map before inputting it into Star module, along with the computational flow involved.
The calculation process of Star block is explained below. The Star block receives an input feature map $x_{in}$ with dimensions $C\times H^{\prime}\times W^{\prime}$, where $C$ is the number of channels, and $H^{\prime}$ and $W^{\prime}$ are the height and width of the feature map, respectively.
  • First, a depthwise convolution is applied to the input feature map $x_{in}$, and $W_{k}$ is th convolution kernel for the k-th channel:
    ${x_{in}}^{\prime}=DWConv(x_{in})=\sum_{k=1}^{C}W_{k}*x_{{in}},$ (11)
  • Next, $1\times 1$ convolution layers are used to transform features, producing two feature maps:
    ${x}_{1}^{\prime}={Conv}_{1\times 1}({x_{in}}^{\prime})=W_{1}\cdot x_{in}^{\prime}+b_{1},$ (12)
    ${x}_{2}^{\prime}={Conv}_{1\times 1}({x_{in}}^{\prime})=W_{2}\cdot x_{in}^{\prime}+b_{2}.$
  • Then, the first feature map undergoes ReLU6 activation, followed by element-wise multiplication with the second feature map ${x_{in}}^{\prime\prime}$, and another another $1\times 1$ convolution layer is applied to generate a new feature map ${x}^{(new)}$:
    ${x_{in}}^{\prime\prime}=min(max(0,{x}_{1}^{\prime}),6)\odot{x}_{2}^{\prime},$ (13)
    ${x}^{(new)}$ $=\text{DWConv}2(\text{Conv}_{1\times 1}(g({x_{in}}^{\prime\prime})))$ (14)
    $=\sum_{k=1}^{C}W_{k}^{new}*g(x_{{in}}^{\prime\prime}).$
  • Finally, a residual connection is established by adding the input feature map ${x_{in}}$ to the newly generated feature map, applying dropout for regularization:
    ${y}={x_{in}}+\text{dropout}({x}^{(new)})$ (15)
II-D LS Detector
The FOCS [18] shows that groupnorm enhances detection head performance in localization and classification. Shared convolutions reduce model parameters, facilitating lightweight design for resource-constrained devices. To address inconsistencies in handling targets of varying scales, we introduce a scale layer for feature adjustment. The specific structure is shown in Fig. 3. These strategies lower parameters and computational load while maintaining accuracy.
Figure 4: The heat-map visualization results feature three rows: the first row illustrates the ground truth, the second row presents the baseline (YOLOv8n), and the third row highlights our proposed method. Red circles represent missed detections.
TABLE I: SPECIFIC EXPERIMENTAL DETAILS
Dataset Epochs Optimizer Batch Lr0 Momentum Weight_decay
SSDD 300 AdamW 16 0.002 0.937 5e-4
HRSID 300 AdamW 16 0.002 0.937 5e-4
TABLE II: ABLATION EXPERIMENTS ON HRSID
WCG WSF LS $\mathrm{mAP_{.50}}(\%)$ $\mathrm{mAP_{.50:.95}}(\%)$ Params(M) FLOPs(G)
× × × 91.1 66.9 3.01 8.1
× × 91.2 66.5 2.43 7.4
× 91.1 65.8 2.13 6.8
91.2 67.6 1.49 5.1
III EXPERIMENT
III-A Datasets
To evaluate the performance of our model, we utilized two publicly available SAR remote sensing datasets: SSDD [19] and HRSID [20]. The SSDD consists of 1160 images, with 928 allocated for training and 232 for testing, as outlined in the original publication. In the case of the HRSID, we adhered to the official partitioning, employing 3642 images for training and 1962 for testing purposes.
III-B Training Details
Our experiments were conducted on a system running Ubuntu 20.04, equipped with an Intel Xeon Silver 4210R CPU and a Nvidia GeForce Tesla V100 16G. And the image size is set to 640. The specific experimental details are shown in Table I.
TABLE III: COMPARISON EXPERIMENTS ON HRSID
Method $\mathrm{mAP_{.50}}(\%)$ $\mathrm{mAP_{.50:.95}}(\%)$ Params(M) FLOPs(G)
Faster R-CNN-FPN 72.0 46.5 41.35 134
RetinaNet 78.9 53.6 36.33 128
YOLOv6n 88.2 62.8 4.23 11.8
YOLOv7-tiny 85.4 57.2 6.01 13.2
YOLOv8n 91.1 66.9 3.01 8.1
YOLOV11n 89.7 65.8 2.59 6.4
Yue et al. [21]* 91.1 66.5 43.42 -
CPoints-Net [22]* 90.5 - 18.64 102.2
FEPS-Net [23]* 90.7 65.7 37.31 -
BL-Net [24]* 88.67 - 47.81 -
FBUA-Net [25]* 90.3 - 36.54 -
$\mathrm{{CS^{n}Net}}$ [26]* 91.2 60.6 12.2 21.7
Ours 91.2 67.6 1.49 5.1
* Denotes data obtained from their papers
TABLE IV: COMPARISON EXPERIMENTS ON SSDD
Method $\mathrm{mAP_{.50}}(\%)$ $\mathrm{mAP_{.50:.95}}(\%)$ Params(M) FLOPs(G)
Faster R-CNN-FPN 94.0 64.0 41.35 134
RetinaNet 90.2 60.0 36.33 128
YOLOv6n 96.9 71.2 4.23 11.8
YOLOv7-tiny 96.4 66.5 6.01 13.2
YOLOv8n 98.2 73.1 3.01 8.1
YOLOV11n 98.0 72.0 2.59 6.4
Yue et al. [21]* 95.7 64.8 43.42 -
CPoints-Net [22]* 96.3 - 18.64 102.2
FEPS-Net [23]* 96.0 73.1 37.31 -
BL-Net [24]* 95.25 - 47.81 -
FBUA-Net [25]* 96.2 - 36.54 -
$\mathrm{{CS^{n}Net}}$ [26]* 97.1 64.9 12.2 21.7
Ours 98.4 72.5 1.49 5.1
* Denotes data obtained from their papers
III-C Evaluation Metrics
To evaluate the proposed model’s performance, we used mean average precision at IoU = 0.5 ($\mathrm{mAP_{.50}}$) and IoU = 0.5:0.05:0.95 ($\mathrm{mAP_{.50:.95}}$), along with the number of parameters (Params) and floating-point operations per second (FLOPs). The mAP metrics assess detection accuracy, while Params and FLOPs indicate parameters and computational complexity.
III-D Ablation Experiments
To validate the module’s effectiveness, we conducted ablation experiments on the HRSID dataset, as shown in Table II. We also performed feature visualization experiments, illustrated in Fig. 4, which demonstrate that our model effectively identifies real targets while reducing the influence of surrounding noise.
III-E Comparison Experiments
We selected several representative object detectors, including the classic two-stage Faster R-CNN-FPN [27], one-stage RetinaNet [18], and YOLOv6n [28], YOLOv7-tiny [29], YOLOv8n and YOLOv11n. Additionally, we compared algorithms specifically designed for target detection in SAR images. The experimental results, presented in Tables III and IV, demonstrate that our method effectively balances parameter quantity and performance, outperforming other methods.
IV CONCLUSIONs
This letter introduces RSNet, a lightweight network for SAR ship detection that enhances accuracy and efficiency through its innovative architecture. The WCG backbone improves performance with fewer parameters, facilitating effective feature extraction and context understanding. The WSF neck optimizes feature integration, while the LS module ensures efficient detection with minimal computational overhead. Future work will explore RSNet’s application to other detection tasks and datasets, further expanding its utility in remote sensing.